OpenStudy (yrivers36):

Write the slope intercept form of a line passing through (-1,4) and (2,-2)

OpenStudy (amistre64):

what do you get when you subtract one point from the other?

OpenStudy (amistre64):

hard to have a conversation when they vanish; but it could be a bad connection ontheir end i spose

OpenStudy (amistre64):

subtract your points and stack y/x for slope: m then pick either point (Px,Py) and construct the equation: y = mx -mPx + Py

OpenStudy (yrivers36):

The answer that I came up is y=3x+1. The teacher says its y=2x+2

OpenStudy (amistre64):

(-1,4) -( 2,-2) ------ -3,6; slope = y/x = 6/-3 = -2 y = -2x +2(-1)+4 = -2x -2+4 = -2x + 2

OpenStudy (amistre64):

youre teachers is closest :)

OpenStudy (yrivers36):

im still unsure. going over your answer

OpenStudy (yrivers36):

isn't the formula y1-y2/x1-x2

OpenStudy (amistre64):

we can test it out to BE sure: y = -2x + 2 ; try out (-1,4) 4 -1 ----------- 4 = 2 + 2 ; thats good y = -2x + 2 ; try out (2,-2) -2 2 ---------- -2 = -4+2 thats good as well

OpenStudy (amistre64):

the formula is just a compact way to write the steps, you still have to do the steps. P(x1 , y1) -Q(x2 , y2) ------------ (x1-x2, y1-y2); slope = y/x = (y1-y2)/(x1-x2) but even that is misleading because it doesnt matter which point you start with

OpenStudy (yrivers36):

m=-2-4[\div-2-1\] and I came up with -6/2 =-3

OpenStudy (amistre64):

your not keeping track of your numbers to well, which happems to me when i try this horizontal conflagerations. I always do a vertical format so that i dont mix up numbers

OpenStudy (amistre64):

if we do "the formula" m = (4--2)/(-1-2) = 6/-3 = -2 or m = (-2-4)/(2--1) = -6/3 = -2

OpenStudy (yrivers36):

ooohhh I see what i did. I didn't change my negative to a positive

OpenStudy (amistre64):

:) good catch

OpenStudy (yrivers36):

oh ok I got it...thank you so much for your help

OpenStudy (amistre64):

youre welcome :)